OpenStudy (anonymous):

Medddallll and Fannnns!!! Consider the vector v=(5,8) What is the angle between v=(5,8) and the positive x-axis Write v in the form (|v|cos theta|v|sin theta)

OpenStudy (amistre64):

\[cos\theta=\frac{\vec u\cdot \vec v}{|\vec u||\vec v|}\]

OpenStudy (amistre64):

or, tan theta = 8/5

OpenStudy (anonymous):

Ok but how would we express the numbers in the equation

OpenStudy (amistre64):

the top is a dot product, the bottom is the product of their lengths

OpenStudy (anonymous):

Ahhh, ok..

OpenStudy (amistre64):

the simplest vector in the posx direction to use is (1,0) its length is 1 and it dots out to the x component of v

OpenStudy (amistre64):

(a,b) (1,0) ----- a+0 = a

OpenStudy (anonymous):

So thats for the vector position.. So what is the equation for the lenths

OpenStudy (amistre64):

since the length of (1,0) is 1 ... and 1|v| = |v| you should be able to determine the length of v by now

OpenStudy (amistre64):

\[cos\theta=\frac{v_x}{|v|}\]

OpenStudy (amistre64):

sin(theta) = sqrt(1-cos^2)

OpenStudy (anonymous):

ok so the length i got was 13... But instead you would want it to be sqrt 13 right

OpenStudy (amistre64):

8^2 + 5^2 = 64+25 = 89, id go with sqrt(89)

OpenStudy (anonymous):

Ok but how would we express that for the angle between v=(5,8) and the positive x axis?

OpenStudy (amistre64):

\[cos\theta=\frac{5}{\sqrt{89}}\] \[\theta=cos^{-1}\left(\frac{5}{\sqrt{89}}\right)\]

OpenStudy (amistre64):

OR\[tan\theta=\frac{8}{5}\] \[\theta=tan^{-1}\left(\frac{8}{5}\right)\]

OpenStudy (anonymous):

Alright, Thank you!

OpenStudy (amistre64):

your welcome

OpenStudy (anonymous):

So then I can plug it in to make it appear on the graph by the vector